If a certain two digit number is divided by sum of its digits the quotient is 7 and remainder is 3. if the digits are interchanged and resulting no is divided by positive difference of digits the quotient is 8 and remainder is 1. find the number
Answer by mananth(16949)

Let x be in ten's place
and y in unit's place
the number is 10x+y
10x+y = (x+y) *7 +3
10x+y = 7x+7v+3
3x-6y = 3
/3
x-2y=1.........................(1)
Digits interchanged , number will be 10y+x
10y+x = 8(x-y)+1
10y+x=8x-8y+1
18y-7x=1.........................(2)
Multiply (1) by 7 and add it to (2)
7x-14y=7
18y-7x=1
4y=8
y=2
Now x-2y=1
plug y
x-4=1
x=5
The number is 52